You are dragging a heavy box with a mass of 90.0 kg across the floor at a constant velocity with a horizontal force of 300.0 N. a) Calculate uk b) A friend tries to assist you by pushing up on the box with a force of 100.0 N (up) . How much less force will you need to exert to keep the box moving at a constant velocity ? 


1
Expert's answer
2021-01-07T10:55:25-0500

a)

"F=u_k mg,"

"u_k=\frac{F}{mg}=\frac{300}{90\cdot 10}=0.33."

b)

"∆F=F-F^{'}=u_k mg-u_k(mg-F_1)=u_k F_1=0.33\cdot 100=33~N."